# The Gross-Koblitz formula and almost circulant matrices related to Jacobi sums

### v3

@@ Abstract.
In this paper, we mainly consider the arithmetic properties of the cyclotomic matrix $B_{p}(k)=\left[J_{p}(\chi^{ki},\chi^{kj})^{-1}\right]_{1\leq i,j\leq(p-1-k)/k}$, where $p$ is an odd prime, $11\leq k<p-$ 1 is a divisor of $p-1$, $\chi$ is a generator of the group of all multiplicative character of the finite field $\mathbb{F}_{p}$ and $J_{p}(\chi^{ki},\chi^{kj})$ is Jacobi sum over $\mathbb{F}_{p}$.
By using the Gross-Koblitz Formula and some $p$-adic tools, we first prove that
@@
where $p-1=kn$. We next establish some theories on almost circulant matrices and then show that
$\det B_{p}(k)=(-1)^{\frac{(n-1)(p+n-1)}{2}}p^{-(n-1)}n^{n-2}a_{p}(k),$
where $U_{k}$ is the set of all $k$-th roots of unity over $\mathbb{F}_{p}$, and $a_{p}(k)$ is the coefficient of $t$ of the minimal polynomial of the algebraic integer $\sum_{y\in U_{k}}(e^{2\pi{\bf i}y/p}-1)$. Also, for $k=1,2$ we obtain explicit expressions of $\det B_{p}(k)$.
@@ 1.2. Background
In 1979 Gross and Koblitz [6] obtained the well-known Gross-Koblitz formula which reveals the relations between the explict values of Gauss sums and the $p$-adic Gamma function (we will give a detailed introduction in Section 2).
@@
Along this line, Chapman [4] and Vsemirnov [15, 16] further investigated some interesting variants of Carlitz’s matrix $C_{p}$. The most well-known result among these variants is Vsemirnov’s ingenious proof on Chapman’s “evil determinant conjecture”. This challenging conjecture states that
@@ 1.4. Main Theorems
(i) Suppose that $a_{p}(k)$ is the coefficient of $t$ of the polynomial $P_{k}(t)$. Then
@@
(i) In Section 5, we will see that $\theta_{k}$ is indeed a generator of the unique intermediate field $L_{n}$ of $\mathbb{Q}(\zeta_{p})$ with $[L_{n}:\mathbb{Q}]=n$, that is, $\mathbb{Q}(\theta_{k})=L_{n}$.
(ii) Although Theorem 1.2(i) states that $\det B_{p}(k)=(-1)^{\frac{(n-1)(p+n-1)}{2}}p^{-(n-1)}n^{n-2}a_{p}(k)$, for an arbitrary divisor $k$ of $p-1$ with $13\leq k<p-$ 1, finding a simple expression of $\det B_{p}(k)$ like (ii) and (iii) in Theorem 1.2 seems difficult. Even for the case of $k=3$, we cannot get a satisfactory expression. The following table lists the values of $a_{p}(3)$ for $p=7,13,19,31,37,43,61,67$.
@@
$\frac{a_{p}(k)}{p}\equiv-\left(\frac{1}{(k-1)!}\right)^{n-2}\frac{1}{(2k)!}\ ({\rm{mod}}\ p).$
@@ 1.5. Outline of This Paper
In Section 2, we will introduce some preliminaries on Gauss sums, Jacobi sums and the $p$-adic Gamma functions. The proof of Theorem 1.1 will be given in Section 3. In Section 4, we will introduce the definition of almost circulant matrices and pose some necessary lemmas. Finally, we will prove Theorem 1.2 in Section 5.
@@
Recall that $p$ is an odd prime. We next introduce the $p$-adic Gamma function $\Gamma_{p}$. For any $n\in\mathbb{Z}^{+}$, we first define
@@
As $\mathbb{Z}^{+}$ is dense in $\mathbb{Z}_{p}$ and $\mathbb{Z}_{p}^{\times}$ is a closed multiplicative group, the $p$-adic Gamma function $\Gamma_{p}:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}^{\times}$ is defined by
@@
We next list some basic properties of $p$-adic Gamma function.
@@
We now state a relationship between the $p$-adic Gamma function and Gauss sums. Let $\pi\in\mathbb{C}_{p}$ be a root of the equation $x^{p-1}+p=0$. Let $\zeta_{\pi}\in\mathbb{C}_{p}$ be a unique primitive $p$-th root of unity such that $\zeta_{\pi}\equiv 1+\pi\ ({\rm{mod}}\ \pi^{2})$. Also, since
@@ 4.1. Circulant Matrices
${\bm{x}}_{l}=\frac{1}{\sqrt{n}}\left(1,\xi,\cdots,\xi^{(n-1)l}\right)^{T}.$
@@
(4.22) $V_{n}E_{n}=E_{n}{\rm diag}\left(\lambda_{0},\lambda_{l},\cdots,\lambda_{n-1}\right).$
@@ 4.2. An Auxiliary Theorem on Almost Circulant Matrices
$\theta_{k}=-k+\sum_{y\in U_{k}}\zeta_{p}^{y}=\sum_{y\in U_{k}}\left(\zeta_{p}-1\right).$
@@
Clearly $\widetilde{D_{p}}(k)$ is the circulant matrix of ${\bm{v}}=(v_{0},v_{1},\cdots,v_{n-1})$ where $v_{0}=-p$ and $v_{i}=G_{p}(\chi^{ki})$ for $i=1,2,\cdots,n-1$. Also, we have
@@
Acknowledgments The authors would like to thank Prof. Hao Pan for his encouragement.
This research was supported by the Natural Science Foundation of China (Grant Nos. 12101321 and 12201291).

### v4

# accepted for publication in Finite Fields and Their Applications

@@ Abstract.
In this paper, we mainly consider arithmetic properties of the cyclotomic matrix $B_{p}(k)=\left[J_{p}(\chi^{ki},\chi^{kj})^{-1}\right]_{1\leq i,j\leq(p-1-k)/k}$, where $p$ is an odd prime, $11\leq k<p-$ 1 is a divisor of $p-1$, $\chi$ is a generator of the group of all multiplicative characters of the finite field $\mathbb{F}_{p}$ and $J_{p}(\chi^{ki},\chi^{kj})$ is Jacobi sum over $\mathbb{F}_{p}$. By using the Gross-Koblitz formula and some $p$-adic tools, we first prove that
@@
where $p-1=kn$. By establishing some theories on almost circulant matrices, we show that
$\det B_{p}(k)=(-1)^{\frac{(n-1)(p+n-1)}{2}}p^{-(n-1)}n^{n-2}a_{p}(k).$
Here $a_{p}(k)$ is the coefficient of $t$ in the minimal polynomial of $\sum_{y\in U_{k}}(e^{2\pi{\bf i}y/p}-1)$, where $U_{k}$ is the set of all $k$-th roots of unity over $\mathbb{F}_{p}$. Also, for $k=1,2$ we obtain explicit expressions of $\det B_{p}(k)$.
@@ 1.2. Background
In 1979 Gross and Koblitz [6] obtained the well-known Gross-Koblitz formula which reveals the relations between the explict values of Gauss sums and the $p$-adic $\Gamma$ function (we will give a detailed introduction in Section 2).
@@
Along this line, Chapman [4] and Vsemirnov [15, 16] further investigated some interesting variants of Carlitz’s matrix $C_{p}$. The most well-known result among these variants is Vsemirnov’s ingenious proof of Chapman’s “evil determinant conjecture”. This challenging conjecture states that
@@ 1.4. Main Theorems
$p^{n-2}\det B_{p}(k)\equiv(-1)^{\frac{(n-1)(p+n-3)}{2}}\left(\frac{1}{k!}\right)^{n-2}\frac{1}{(2k)!}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
@@
(i) Suppose that $a_{p}(k)$ is the coefficient of $t$ in the polynomial $P_{k}(t)$. Then
@@
(i) Note that
$(-1)^{\frac{p+1}{2}+\lfloor\frac{p-3}{4}\rfloor}=\left(\frac{-2}{p}\right).$
(ii) In Section 5, we will see that $\theta_{k}$ is indeed a generator of the unique intermediate field $L_{n}$ of $\mathbb{Q}(\zeta_{p})$ with $[L_{n}:\mathbb{Q}]=n$, that is, $\mathbb{Q}(\theta_{k})=L_{n}$.
(iii) Although Theorem 1.2(i) states that $\det B_{p}(k)=(-1)^{\frac{(n-1)(p+n-1)}{2}}p^{-(n-1)}n^{n-2}a_{p}(k)$, for an arbitrary divisor $k$ of $p-1$ with $13\leq k<p-$ 1, finding a simple expression of $\det B_{p}(k)$ like (ii) and (iii) in Theorem 1.2 seems difficult. Even for the case of $k=3$, we cannot get a satisfactory expression. The following table lists the values of $a_{p}(3)$ for $p=7,13,19,31,37,43,61,67$.
@@
$\frac{a_{p}(k)}{p}\equiv-\left(\frac{1}{(k-1)!}\right)^{n-2}\frac{1}{(2k)!}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
@@ 1.5. Outline of This Paper
In Section 2, we will introduce some preliminaries on Gauss sums, Jacobi sums and the $p$-adic $\Gamma$ functions. The proof of Theorem 1.1 will be given in Section 3. In Section 4, we will introduce the definition of almost circulant matrices and pose some necessary lemmas. Finally, we will prove Theorem 1.2 in Section 5.
@@
Recall that $p$ is an odd prime. We next introduce the $p$-adic $\Gamma$ function $\Gamma_{p}$. For any $n\in\mathbb{Z}^{+}$, we first define
@@
As $\mathbb{Z}^{+}$ is dense in $\mathbb{Z}_{p}$ and $\mathbb{Z}_{p}^{\times}$ is a closed multiplicative group, the $p$-adic $\Gamma$ function $\Gamma_{p}:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}^{\times}$ is defined by
@@
We next list some basic properties of $p$-adic $\Gamma$ function.
@@
We now state a relationship between the $p$-adic $\Gamma$ function and Gauss sums. Let $\pi\in\mathbb{C}_{p}$ be a root of the equation $x^{p-1}+p=0$. Let $\zeta_{\pi}\in\mathbb{C}_{p}$ be the unique primitive $p$-th root of unity such that $\zeta_{\pi}\equiv 1+\pi\ ({\rm{mod}}\ \pi^{2})$. Also, since
@@ 4.1. Circulant Matrices
${\bm{x}}_{l}=\frac{1}{\sqrt{n}}\left(1,\xi^{l},\cdots,\xi^{(n-1)l}\right)^{T}.$
@@
(4.3) $V_{n}E_{n}=E_{n}{\rm diag}\left(\lambda_{0},\lambda_{1},\cdots,\lambda_{n-1}\right).$
@@ 4.2. An Auxiliary Theorem on Almost Circulant Matrices
$\theta_{k}=-k+\sum_{y\in U_{k}}\zeta_{p}^{y}=\sum_{y\in U_{k}}\left(\zeta^{y}_{p}-1\right).$
@@
Clearly $\widetilde{D_{p}}(k)$ is the transpose of the circulant matrix of ${\bm{v}}=(v_{0},v_{1},\cdots,v_{n-1})$ where $v_{0}=-p$ and $v_{i}=G_{p}(\chi^{ki})$ for $i=1,2,\cdots,n-1$. Also, we have
@@
Declaration of competing interest The authors declare that they have no conflict of interest.
Data availability No data was used for the research described in the article.
Acknowledgments The authors would like to thank two referees for their helpful comments. The authors also thank Professor Hao Pan for his encouragement.
We dedicate this paper to our advisor Professor Zhi-Wei Sun on the occasion of his 60th birthday.
This research was supported by the Natural Science Foundation of China (Grant Nos. 12101321 and 12201291). The first author was also supported by Natural Science Foundation of Nanjing University of Posts and Telecommunications (Grant No. NY224107).